A country with a civilian population of 120, 000 (all over age 16) has 100,000 employed and 10, 000 unemployed persons, of which 5, 000 are frictionally unemployed and another 3, 000 structurally unemployed. Given the above data, complete the following questions. Show the work.
1. What is the size of the labor force?
2. What is the actual unemployment rate?
What is the labor force participation rate?
4. What is the employment-population ratio?
5. What is the natural rate of unemployment?
1
Expert's answer
2020-11-09T06:59:36-0500

1) "labor force = unemployed+ employed"

=100000+ 10000=110000

2) unemployment rate ,"U=(unemployed\/ labor)\u00d7100"

U=(100000/110000)×100= 9.09%

3) labor participation, "P=(labor\/population)\u00d7100"

P=(110000/120000)×100

P=91.7%

4) employment: population

=100000/120000

=1:1

5) natural rate,

"N=(( frictional+structural)\/labor )\u00d7100"

N=(7000/110000)×100= 6.4%